# biblio.bib

### v2

@@ Appendix B New real solutions for $n=6$
In this appendix we give cartesian coordinates on $S^{3}$ for the new real solutions of six points, as described in Section 11.6.
@@
Based on the geometrical interpretation of Section 12, and inspired by the Föppl notation, we propose the following notation for these configurations:
(Real $\begin{array}[]{ll}\text{({Real 1})}\quad 1:{4}_{S^{2}}:1,&\text{({Real 2})}\quad{4_{\text{Eq.}}:2_{S^{2}}}\\ \text{({Real 3})}\quad 1:5_{S^{2}},&\text{({Real 4})}\quad 3_{S^{1}}\times 3_{S^{1}};\end{array}$
where the last one uses $\times$ since the configuration lies in $S^{1}\times S^{1}$.
@@ Appendix C Projected Hessian eigenvalues
The eigenvalues of the projected Hessian are given in Tables 12, 13 and 14. The associated code can be found in the SymPy notebook “classifyHessian”. For 1:5 and 3:3 we could not calculate the exact eigenvalues, and we report the eigenvalues calculated with floating point arithmetic. To prove that these configurations are saddle, we give feasible directions where the quadratic form of the Hessian is negative. To find a negative direction for 1:5, we move the first two points of the roots of unity, one upward and the other downward. The associated value of the quadratic form of the Hessian is:

### v3

@@ Appendix B New real solutions for $n=6$
In this appendix we give cartesian coordinates on $S^{3}$ for the new real solutions of six points, as described in Section 4.2.6.
@@
Based on the geometrical interpretation of Section 4.3, and inspired by the Föppl notation, we propose the following notation for these configurations:
$\begin{array}[]{ll}\text{({Real 1})}\qquad 1:{4}_{S^{2}}:1,&\text{({Real 2})}\qquad{4_{\text{Eq.}}:2_{S^{2}}}\\ \text{({Real 3})}\qquad 1:5_{S^{2}},&\text{({Real 4})}\qquad 3_{S^{1}}\times 3_{S^{1}}.\end{array}$
@@ Appendix C Projected Hessian eigenvalues
The eigenvalues of the projected Hessian are given in Tables 18, 19 and 20. The associated code can be found in the SymPy notebook “classifyHessian”. For 1:5 and 3:3 we could not calculate the exact eigenvalues, and we report the eigenvalues calculated with floating point arithmetic. To prove that these configurations are saddle, we give feasible directions where the quadratic form of the Hessian is negative. To find a negative direction for 1:5, we move the first two points of the roots of unity, one upward and the other downward. The associated value of the quadratic form of the Hessian is: